OpenStudy (anonymous):

factorize this expression; X'3 - 1

OpenStudy (anonymous):

\[x ^{3}-1\]

OpenStudy (akshay_budhkar):

x^3-1=(x-1)(x^2+x+1)

OpenStudy (jwt625):

(x-1)(x^2+x+1)

OpenStudy (akshay_budhkar):

you can usr the formula a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (akshay_budhkar):

get it?